A certain dog whistle has a frequency of 35.10 kHz.  A person blows the whistle while riding in the back 

of a sweet Sebring convertible.  With what minimum velocity must the driver go in order for a person standing on the side of the road to hear the whistle as the car drives away? 

[recall that humans can only hear sounds up to 20,000 Hz]



1
Expert's answer
2021-01-25T15:59:19-0500

The Doppler's effect equation gives

"f-f_0=-f_0 \frac{v}{c}"

"v=c\times \frac{f_0-f}{f_0}"

"v=\rm 340\: m/s\times \frac{35100-20000}{35100}=146\: m/s"